# Log Neural Controlled Differential Equations: The Lie Brackets Make a Difference

### v2

@@ 4.2 Toy Dataset
On this dataset, all models use a hidden state of dimension $64$ and Adam with a learning rate of $0.0001$ (Kingma & Ba, 2017). LRU, S5, S6, and MAMBA use six blocks. NRDE and Log-NCDE take $r_{i+1}-r_{i}$ to be $4$ observations and the signature truncation depth $N$ to be $2$. Full hyperparameter choices are given in Appendix C.3.
@@ 5.1 Toy Dataset
Figure 5 compares the performance of the models considered in this paper on the four different toy dataset classifications. As expected, given that the classifications considered are solutions to CDEs, NCDE’s are the best performing model. Since NRDEs and Log-NCDEs are fixed to $r_{i+1}-r_{i}$ being $4$ observations and $N=2$, they are both approximations of a CDE. Notably, Log-NCDEs consistently outperform NRDEs, providing empirical evidence that NRDEs do not always accurately learn the Lie bracket structure of $\bar{f}_{\theta}$. From the stacked recurrent models, MAMBA achieves the highest validation accuracy on all four datasets. However, the performance of all the stacked recurrent models decreases drastically as the number of channels the label depends on increases.
@@
EigenWorms $\underline{\mathbf{85.0\pm 6.2}}$ $83.9\pm 4.1$ $\underline{\mathbf{85.0\pm 16.1}}$ $70.9\pm 15.8$ $62.2\pm 2.2$ $77.2\pm 7.1$ $82.8\pm 2.7$
EthanolConcentration $23.8\pm 2.8$ $25.6\pm 3.5$ $26.4\pm 6.4$ $27.9\pm 4.5$ $22.0\pm 1.0$ $\underline{31.4\pm 4.5}$ $\mathbf{35.9\pm 6.1}$
Heartbeat $\mathbf{78.1\pm 7.6}$ $73.9\pm 3.1$ $\underline{76.5\pm 8.3}$ $76.2\pm 3.8$ $68.1\pm 5.8$ $73.9\pm 2.6$ $74.2\pm 2.0$
MotorImagery $51.9\pm 8.6$ $53.0\pm 3.9$ $51.3\pm 4.7$ $47.7\pm 4.5$ $51.6\pm 6.2$ $\underline{54.0\pm 7.8}$ $\mathbf{57.2\pm 5.6}$
SelfRegulationSCP1 $\underline{84.5\pm 4.6}$ $\mathbf{87.1\pm 2.1}$ $82.8\pm 2.7$ $80.7\pm 1.4$ $80.0\pm 2.0$ $76.7\pm 5.6$ $82.1\pm 1.4$
SelfRegulationSCP2 $47.4\pm 4.0$ $\mathbf{55.1\pm 3.3}$ $49.9\pm 9.4$ $48.2\pm 3.9$ $49.1\pm 6.2$ $48.1\pm 11.4$ $\underline{54.0\pm 2.6}$
Av. $61.8$ $\underline{63.1}$ $62.0$ $58.6$ $55.5$ $60.2$ $\mathbf{64.4}$
Av. Rank $3.6$ $\underline{3.1}$ $3.3$ $4.8$ $6.0$ $4.6$ $\mathbf{2.7}$
@@
LRU 5484.3 379.5
S5 5327.7 441.4
S6 2608.0 578.1
MAMBA 4450.3 1609.1
NCDE 1674.3 7720.7
NRDE 3338.3 8338.3
Log-NCDE 2661.3 2200.1
@@ 5.2 UEA-MTSCA
Table 1 reports the average and standard deviation of each model’s test set accuracy over five different splits of the data. On the datasets considered, NCDEs achieve the lowest average accuracy. NRDEs improve upon NCDEs in both average accuracy and rank, but are still outperformed by three of the stacked recurrent models, LRU, S5, and S6. Log-NCDEs are the best performing model on average test set accuracy and average rank. Compared to NRDEs, they achieve a higher average test set accuracy on all six datasets and a lower standard deviation on five datasets. These results highlight the improvement in performance which can be achieved by calculating the Lie brackets.
@@ 5.3 PPG-DaLiA
Table 3 contains the average and standard deviation of each model’s test set mean squared error on the PPG-DaLiA dataset. In contrast to the UEA-MTSCA experiments, MAMBA is the best performing stacked recurrent model and NRDEs do not improve over NCDEs. However, Log-NCDEs still achieve the best performance, obtaining the lowest average test set mean squared error and the second lowest standard deviation.
@@
MAMBA $\underline{10.65\pm 2.20}$
NCDE $13.77\pm 0.41$
NRDE $18.79\pm 2.69$
Log-NCDE $\mathbf{9.20\pm 0.55}$
@@ C.3 Toy Dataset Details
On the toy dataset, all models use a hidden state of dimension $64$ and Adam with a learning rate of $0.0001$ (Kingma & Ba, 2017). LRU, S5, S6, and MAMBA use $6$ blocks and S5, S6, and MAMBA use a state dimension of $64$. S5 uses $2$ initialisation blocks and MAMBA uses a convolution dimension of $4$ and an expansion factor of $2$. NCDEs, NRDEs, and Log-NCDEs use a FCNN with width $128$ and depth $3$ as their vector field. Furthermore, they all use Heun as their differential equation solver with a fixed stepsize of $0.01$. NRDEs and Log-NCDEs use a Log-ODE step of $4$ and a signature truncation depth of $2$. Log-NCDEs do not use any $\mathrm{Lip}(\gamma)$ regularisation, i.e. $\lambda=0$.
@@ C.4 UEA-MTSCA and PPG-DaLiA Details
  Learning Rate $10^{-3}$ EW, MI, SCP2, PPG EW, EC, PPG EW, HB, MI, SCP2 EW, EC, PPG
$10^{-4}$ EC, HB, SCP1 HB, MI, SCP1, SCP2 SCP1, PPG HB, SCP2
$10^{-5}$ EC MI, SCP1
  Include Time True EW, SCP1, PPG EW, EC, MI, SCP2, PPG EC, HB, MI, PPG EW, EC, SCP2
False EC, HB, MI, SCP2 HB, SCP1 EW, SCP1, SCP2 HB, MI, SCP1, PPG
  Hidden Dimension 16 PPG EC, SCP1, PPG EW, EC, HB, MI, SCP2 EW
64 MI, SCP1 HB, SCP2 SCP1, PPG EC, HB, SCP2
128 EW, EC, HB, SCP2 EW, MI MI, SCP1, PPG
  Number of Layers 2 EC EW, HB, SCP2 SCP1, SCP2, PPG MI, SCP1
4 MI, SCP1 EW, EC, HB, MI EC, HB, PPG
6 EW, HB, MI, SCP1, SCP2, PPG EC, PPG EW, SCP2
  State Dimension 16 EW, EC EC, HB EC, HB, SCP1 SCP1
64 SCP2 SCP1 EW, PPG EW, MI, SCP2, PPG
256 HB, MI, SCP1, PPG EW, MI, SCP2, PPG MI, SCP2 EC, HB
  S5 Initialisation Blocks 2 EC
@@ 4 EC, HB
  Learning Rate $10^{-3}$ HB, PPG EW, SCP1, SCP2, PPG HB, SCP1, SCP2, PPG
$10^{-4}$ EW, MI, SCP1 EC, HB, MI EW, EC, MI
$10^{-5}$ EC, SCP2
  Include Time True EW, EC, SCP2, PPG EW, PPG EC, HB, MI, SCP2, PPG
False HB, MI, SCP1 EC, HB, MI, SCP1, SCP2 EW, SCP1
  Hidden Dimension 16 EC, HB, SCP1, PPG EW, SCP1, SCP2, PPG HB
64 SCP2 EW, EC, MI, SCP2, PPG
128 EW, MI EC, HB, MI SCP1
  Vector Field (Depth, Width) (2, 32) HB SCP1 HB
(3, 64) EC SCP1
(3, 128) EW, SCP1, PPG EC, SCP2 EC, MI, SCP2
(4, 128) MI, SCP2 EW, HB, MI, PPG EW, PPG
  Log-ODE (Depth, Step) (1, 1) EC, MI, SCP2
(2, 2) HB EC, HB
(2, 4) SCP1 MI
(2, 8) SCP2
(2, 12) EW, SCP1
(2, 16) EW
(1, 10) PPG
@@ C.5 Additional Memory and run time Results
Although the time per training step is lower for stacked recurrent models than NCDEs, NRDEs, or Log-NCDEs, they also require more training steps to converge. Additionally, NCDEs, NRDEs, and Log-NCDEs require less GPU memory. The largest contributors to the average run time of NCDEs are the datasets with the most observations, EigenWorms and MotorImagery. The positive impact of the Log-ODE method on computational burden is demonstrated empirically by the decrease in run time achieved by NRDEs and Log-NCDEs on EigenWorms when using a depth22-2- 2 Log-ODE method. When a depth11-1- 1 Log-ODE method is used, such as NRDEs on MotorImagery, the same decrease is not observed. Section 3.4 demonstrated that Log-NCDEs and NRDEs have the same asymptotic computational complexity. However, when using a depth22-2- 2 Log-ODE approximation and the same stepsize, NRDEs and Log-NCDEs exhibit drastically different run times on Heartbeat, a high-dimensional dataset. This difference is partly explained by the model’s having different optimal hyperparameter choices, but even when using identical hyperparameters to the NRDE, Log-NCDE’s run time for $1000$ steps of training is $3017$ seconds, whereas NRDE’s run time is $20384$ seconds. The remaining difference in run time is due to calculating the JVPs of $f_{\theta}$ using a batched function, as discussed in Section 3.4. If instead the JVPs are calculated using a for-loop, then Log-NCDEs run time increases to $15551$ seconds.
@@
Figure 6: Memory, time for $1000$ steps, number of steps, and approximate total time for each model and dataest from the UEA-MTSCA considered in this paper on an NVIDIA RTX 4090. The following abbreviations are used: EigenWorms (EW), EthanolConcentration (EC), Heartbeat (HB), MotorImagery (MI), SelfRegulationSCP1 (SCP1), and SelfRegulationSCP2 (SCP2).

### v3

@@ 4.2 Toy Dataset
On this dataset, all models use a hidden state of dimension $64$ and Adam with a learning rate of $0.0003$ (Kingma & Ba, 2017). LRU, S5, S6, and MAMBA use six blocks. NRDE and Log-NCDE take $r_{i+1}-r_{i}$ to be $4$ observations and the signature truncation depth $N$ to be $2$. Full hyperparameter choices are given in Appendix C.3.
@@ 5.1 Toy Dataset
Figure 5 compares the performance of the models considered in this paper on the four different toy dataset classifications. As expected, given that the classifications considered are solutions to CDEs, NCDE’s are the best performing model. Since NRDEs and Log-NCDEs are fixed to $r_{i+1}-r_{i}$ being $4$ observations and $N=2$, they are both approximations of a CDE. Notably, Log-NCDEs consistently outperform NRDEs, providing empirical evidence that NRDEs do not always accurately learn the Lie bracket structure of $\bar{f}_{\theta}$. All of the stacked recurrent models perform well when the label depends on one or two channels. However, their performance begins to decrease for three channels, and only MAMBA performs well when the label depends on four channels.
@@
EigenWorms $\mathbf{87.8\pm 2.8}$ $81.1\pm 3.7$ $85.0\pm 16.1$ $70.9\pm 15.8$ $75.0\pm 3.9$ $83.9\pm 7.3$ $\underline{85.6\pm 5.1}$
EthanolConcentration $21.5\pm 2.1$ $24.1\pm 4.3$ $26.4\pm 6.4$ $27.9\pm 4.5$ $\underline{29.9\pm 6.5}$ $25.3\pm 1.8$ $\mathbf{34.4\pm 6.4}$
Heartbeat $\mathbf{78.4\pm 6.7}$ $\underline{77.7\pm 5.5}$ $76.5\pm 8.3$ $76.2\pm 3.8$ $73.9\pm 2.6$ $72.9\pm 4.8$ $75.2\pm 4.6$
MotorImagery $48.4\pm 5.0$ $47.7\pm 5.5$ $\underline{51.3\pm 4.7}$ $47.7\pm 4.5$ $49.5\pm 2.8$ $47.0\pm 5.7$ $\mathbf{53.7\pm 5.3}$
SelfRegulationSCP1 $82.6\pm 3.4$ $\mathbf{89.9\pm 4.6}$ $82.8\pm 2.7$ $80.7\pm 1.4$ $79.8\pm 5.6$ $80.9\pm 2.5$ $\underline{83.1\pm 2.8}$
SelfRegulationSCP2 $51.2\pm 3.6$ $50.5\pm 2.6$ $49.9\pm 9.5$ $48.2\pm 3.9$ $53.0\pm 2.8$ $\underline{\mathbf{53.7\pm 6.9}}$ $\underline{\mathbf{53.7\pm 4.1}}$
Av. $61.7$ $61.8$ $\underline{62.0}$ $58.6$ $60.2$ $60.6$ $\mathbf{64.3}$
Av. Rank $\underline{3.5}$ $4.0$ $\underline{3.5}$ $5.5$ $4.5$ $4.9$ $\mathbf{2.1}$
@@
LRU 4121.67 466.09
S5 2815.00 244.78
S6 2608.00 578.15
MAMBA 4450.33 1553.83
NCDE 1759.67 6649.91
NRDE 2676.33 7284.20
Log-NCDE 1999.67 2128.32
@@ 5.2 UEA-MTSCA
Table 1 reports the average and standard deviation of each model’s test set accuracy over five data splits. MAMBA achieves the lowest average accuracy. NCDEs and NRDEs have similar accuracies except on EigenWorms, the dataset with the most observations, where NRDEs significantly improve over NCDEs. However, NRDEs are still outperformed on average accuracy by three stacked recurrent models: LRU, S5, and S6. Log-NCDEs are the best performing model on average accuracy and rank. Compared to NRDEs, they achieve an equal or higher average accuracy on all six datasets and a lower standard deviation on four datasets. These results highlight the improvement in performance which can be achieved by calculating the Lie brackets.
@@ 5.3 PPG-DaLiA
Table 3 contains the average and standard deviation of each model’s test set mean squared error on the PPG-DaLiA dataset. In contrast to the UEA-MTSCA experiments, MAMBA is the best performing stacked recurrent model. However, Log-NCDEs still achieve the best performance, obtaining the lowest average test set mean squared error and the second lowest standard deviation.
@@
MAMBA $10.65\pm 2.20$
NCDE $13.54\pm 0.69$
NRDE $\underline{9.90\pm 0.97}$
Log-NCDE $\mathbf{9.56\pm 0.59}$
@@ C.3 Toy Dataset Details
On the toy dataset, all models use a hidden state of dimension $64$ and Adam with a learning rate of $0.0003$ (Kingma & Ba, 2017). LRU, S5, S6, and MAMBA use $6$ blocks and S5, S6, and MAMBA use a state dimension of $64$. S5 uses $2$ initialisation blocks and MAMBA uses a convolution dimension of $4$ and an expansion factor of $2$. NCDEs, NRDEs, and Log-NCDEs use a FCNN with width $128$ and depth $3$ as their vector field. Furthermore, they all use Heun as their differential equation solver with a fixed stepsize of $0.01$. NRDEs and Log-NCDEs use a Log-ODE step of $4$ and a signature truncation depth of $2$. Log-NCDEs do not use any $\mathrm{Lip}(\gamma)$ regularisation, i.e. $\lambda=0$.
@@ C.4 UEA-MTSCA and PPG-DaLiA Details
  Learning Rate $10^{-3}$ EW, MI, SCP1, SCP2, PPG HB, MI, SCP1, PPG EW, HB, MI, SCP2 EW, EC, PPG
$10^{-4}$ HB EW, SCP2 SCP1, PPG HB, SCP2
$10^{-5}$ EC EC EC MI, SCP1
  Include Time True EC, HB, SCP2 EW, EC, SCP2, PPG EC, HB, MI, PPG EW, EC, SCP2
False EW, MI, SCP1, PPG HB, MI, SCP1 EW, SCP1, SCP2 HB, MI, SCP1, PPG
  Hidden Dimension 16 MI MI, SCP2, PPG EW, EC, HB, MI, SCP2 EW
64 EW, EC, SCP1, SCP2 EW SCP1, PPG EC, HB, SCP2
128 HB, PPG EC, HB, SCP1 MI, SCP1, PPG
  Number of Layers 2 HB, SCP1, SCP2 EW, EC, SCP2 SCP1, SCP2, PPG MI, SCP1
4 EW HB EW, EC, HB, MI EC, HB, PPG
6 EC, MI, PPG MI, SCP1, PPG EW, SCP2
  State Dimension 16 EC, SCP1, SCP2 EW, EC, HB, SCP1 EC, HB, SCP1 SCP1
64 EW MI, SCP2, PPG EW, PPG EW, MI, SCP2, PPG
256 HB, MI, PPG MI, SCP2 EC, HB
  S5 Initialisation Blocks 2 SCP2, PPG
@@ 4 EC, HB
  Learning Rate $10^{-3}$ EW, EC, HB, MI, SCP2, PPG EW, EC, HB, SCP1, PPG EW, HB, MI, PPG
$10^{-4}$ SCP1 MI, SCP2 EC, SCP1, SCP2
$10^{-5}$
  Include Time True EW, EC, HB, MI, PPG EC, SCP1, SCP2, PPG EW, EC, HB, PPG
False SCP1, SCP2 EW, HB, MI MI, SCP1, SCP2
  Hidden Dimension 16 MI, PPG HB, MI
64 EW, EC, HB, SCP1 EC, SCP1
128 EW, EC, HB, SCP1, SCP2 MI, SCP2, PPG EW, SCP2, PPG
  Vector Field (Depth, Width) (2, 32) EW, SCP2
(3, 64) EW EC, MI, PPG
(3, 128) EW HB HB, SCP1
(4, 128) EC, HB, MI, SCP1, SCP2, PPG EC, MI, SCP1, SCP2, PPG
  Log-ODE (Depth, Step) (1, 1) EC, MI, SCP1, SCP2 EC
(2, 2) HB HB
(2, 4) EW SCP2
(2, 8)
(2, 12) EW
(2, 16) MI, SCP1
(1, 10) PPG PPG
@@ C.5 Additional Memory and run time Results
Although the time per training step is lower for stacked recurrent models than NCDEs, NRDEs, or Log-NCDEs, they also require more training steps to converge. Additionally, NCDEs, NRDEs, and Log-NCDEs require less GPU memory. The largest contributors to the average run time of NCDEs are the datasets with the most observations, EigenWorms and MotorImagery. The positive impact of the Log-ODE method on computational burden is demonstrated empirically by the decrease in run time achieved by NRDEs and Log-NCDEs on EigenWorms when using a depth22-2- 2 Log-ODE method. When a depth11-1- 1 Log-ODE method is used, such as NRDEs on MotorImagery, the same decrease is not observed. Section 3.4 demonstrated that Log-NCDEs and NRDEs have the same asymptotic computational complexity. However, when using a depth22-2- 2 Log-ODE approximation and the same stepsize, NRDEs and Log-NCDEs exhibit drastically different run times on Heartbeat, a high-dimensional dataset. This difference is partly explained by the model’s having different optimal hyperparameter choices, but even when using identical hyperparameters to the NRDE, Log-NCDE’s run time for $1000$ steps of training is $1673$ seconds, whereas NRDE’s run time is $9539$ seconds. The remaining difference in run time is due to calculating the JVPs of $f_{\theta}$ using a batched function, as discussed in Section 3.4. If instead the JVPs are calculated using a for-loop, then Log-NCDEs run time increases to $17045$ seconds.
@@
Figure 6: Memory, time for $1000$ steps, number of steps, and approximate total time for each model and dataeset from the UEA-MTSCA considered in this paper on an NVIDIA RTX 4090. The following abbreviations are used: EigenWorms (EW), EthanolConcentration (EC), Heartbeat (HB), MotorImagery (MI), SelfRegulationSCP1 (SCP1), and SelfRegulationSCP2 (SCP2).